OpenStudy (anonymous):

Determine if the function f is an exponential function. If so, identify the base. If not, why not? f(x) = x^(-3.141)

OpenStudy (anonymous):

@BangkokGarrett I have three more questions. Could you help me. I believe this one is not an exponential function?

OpenStudy (anonymous):

No. The base is not a constant. It's a variable

OpenStudy (anonymous):

Find the y-axis intercept for the function: f(x) = -e^(3x+1) - 5. (0, -5) (-e-5, 0) (-5, 0) (0, -e-5)

OpenStudy (anonymous):

To find a y-intercept, set x=0 and find out what y (or f(x) in your case) equals

OpenStudy (anonymous):

Would it be (0, -e-5)?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

For the graph below, which of the following is a possible function for g? g(x) = (2/11)^x g(x) = (1/4)^x g(x) = 3^(-2x) g(x) = (2/3)^x Give me a second for the graph

OpenStudy (anonymous):

Look at what happens when you plug in -1 to each of those functions 1st one: about 5.5 2nd: about 4 3rd: 9 4th: 1.5 On your graph, what is the value of the blue line when x = -1. It's hard to tell 'cause the y-axis is only marked at 1, but I think I can tell which one it probably is.

OpenStudy (anonymous):

Its either the first or second one. I think its the second one?

OpenStudy (anonymous):

yeah, me too.

OpenStudy (anonymous):

Thank you so much for your help today!

OpenStudy (anonymous):

No problem. Are you in England by the way?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

oh. ok. just wondering, because i saw your file was named "maths.pdf". They say "maths" a lot in England.

OpenStudy (anonymous):

oh lol. No I already had a file named math so i just added an s to it.

OpenStudy (anonymous):

I see. Good luck with your maths!